OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -35 ? -89 -86 -83 -80

OpenStudy (whpalmer4):

Do you know the formula for computing an arbitrary term in an arithmetic sequence?

OpenStudy (anonymous):

Is it an=a1+(n+1)d

OpenStudy (whpalmer4):

Very close... \[a_n = a_1 + (n-1)d\] As an aid to remembering how the formula is written, remember that \[a_1 = a_1\]so\[a_n = a_1+(n-1)d\]and when \(n=1\) then we have \[a_1 = a_1 + (n-1)d\]\[a_1 = a_1 + (0)d\]\[a_1=a_1\] If you had a plus sign in the \((n-1)\), to make that part equal 0 we would have to be talking about \(n=-1\) but we don't do that...

OpenStudy (whpalmer4):

Okay, so we know that \[a_n = a_1 + (n-1)d\]\[a_1 = 13\]\[a_{17} = -35\] Can you use that information to find \(d\)?

OpenStudy (anonymous):

I've calculated it to be three, but this is where I believe I have already messed up

OpenStudy (whpalmer4):

well, let's check: \[a_n = a_1 + (n-1)(3)\]\[a_{17} = -35 = 13 + (17-1)3\]\[-35 = 13+(16)(3)\]Yeah, that isn't going to work out. But we can set it up the other way and work back to \(d\): \[-35 = 13 + (17-1)d\] What do you get if you solve that for \(d\)?

OpenStudy (anonymous):

-3?

OpenStudy (whpalmer4):

Very good. \[-35 -13 = 13-13 + 16d\]\[-48 = 16d\]\[-3=d\] So what is our formula for \(a_n\) in this sequence, and what will it give us for the value of \(a_{35}=\) if we ask nicely? :-)

OpenStudy (anonymous):

Would it be a35 = 13 + (17-1)(-3)?

OpenStudy (whpalmer4):

write out the general formula for \(a_n\) first...

OpenStudy (anonymous):

an =a1 + (n-1) d?

OpenStudy (whpalmer4):

right, now plug in the known values for \(a_1\) and \(d\)

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

an = 13 + (n-1)(-3)

OpenStudy (whpalmer4):

Yes. Now find the value of \(a_{35}\) with that formula. Hint: \(n\ne 17\)

OpenStudy (anonymous):

I got -6.33333... I know this can't be right :c

OpenStudy (whpalmer4):

just put the numbers in the formula and write it out for me...

OpenStudy (anonymous):

a35 = 13 + (n-1) (-3)

OpenStudy (whpalmer4):

I'm glad you recognize that that cannot be the right answer — way too many students lack that sense! what is the value of \(n\)?

OpenStudy (anonymous):

Thank you c: Is n 35?

OpenStudy (whpalmer4):

yes. it's the same number as the subscript of \(a_n\)... (I agree that is a somewhat confusing point at first!)

OpenStudy (anonymous):

So I solve a35 = 13 + (35 - 1) (-3)?

OpenStudy (anonymous):

@whpalmer4? :)

OpenStudy (whpalmer4):

that would be the one...

OpenStudy (anonymous):

-89? :D

OpenStudy (whpalmer4):

you got it!

OpenStudy (anonymous):

Thank you for your time and help! I appreciate it very much :)

OpenStudy (whpalmer4):

so here's a related problem, of a sort I'm confident you'll get sooner or later

OpenStudy (whpalmer4):

might as well learn how to do it now, right? :-)

OpenStudy (anonymous):

sure!

OpenStudy (whpalmer4):

say you have a sequence where you know \[a_3=5\]and\[a_7 = -3\]Can you find the value of \(a_{19}\)?

OpenStudy (anonymous):

Would I solve like here? Or sorta similar at least, this problems looks kind of like the one we just solved. :)

OpenStudy (whpalmer4):

well, the wrinkle is that you don't know the value of \(a_1\)...

OpenStudy (anonymous):

Oh alright, I haven't learned this yet in class. I'd be happy to watch how to solve though :)

OpenStudy (whpalmer4):

but that turns out not to matter, so long as you know two different values of the sequence

OpenStudy (anonymous):

Okay

OpenStudy (whpalmer4):

why don't you write out the equations (with the numbers plugged in) for both values of the sequence: \(a_3 = 5\) and \(a_7 = -3\) that may give you a clue as to how to solve it...

OpenStudy (anonymous):

Well, I don't really have time to solve it myself, I'm going to a family get-together here soon, but I can watch if we do it quickly

OpenStudy (anonymous):

I'd be happy to learn :)

OpenStudy (whpalmer4):

oh, you'll be surprised at how quickly you can do it. just write out the formula, just like you did for the previous problem. I'll do one, you do one. \[a_3 = 5 = a_1 + (5-1)d\] \[5 = a_1 + 4d\]

OpenStudy (whpalmer4):

can you do the one for \(a_7 = -3\)?

OpenStudy (whpalmer4):

\[a_7 = -3 = a_1 + (7-1)d\]\[-3=a_1 + 6d\] Now we have a system of two equations in two unknowns: \[5 = a_1 + 4d\]\[-3 = a_1 + 6d\] You've solved things like that before, right?

OpenStudy (whpalmer4):

Burp, brain fart! \[a_3 = 5 = a_1 + (3-1)d\]\[5 = a_1 + 2d\]\[a_7 = -3 = a_1 + (7-1)d\]\[-3 = a_1+6d\] now that's a system of two equations in two unknowns: \[5 = a_1 + 2d\]\[-3 = a_1 + 6d\]multiply first equation by -1 and add them together: \[-1*5 + -3 = -1*a_1 + a_1 + -1*2d + 6d\]\[-8 = 4d\]\[d=-2\] Now we plug \(d = -2\) into either of the equations to find \(a_1\): \[a_3 = a_1 + (n-1)(-2)\]\[5 = a_1 + 2(-2)\]\[5 = a_1 -4\]\[a_1 = 9\] So our sequence is \[a_n = 9 + (n-1)(-2)\] and \[a_{19} = 9 + (19-1)(-2) = 9 -36 = -27\]

OpenStudy (whpalmer4):

Here's the first 19 terms of that sequence, so you can check my work :-) \[\begin{array}{cc} n & a_n \\ \hline \\1 & 9 \\ 2 & 7 \\ 3 & 5 \\ 4 & 3 \\ 5 & 1 \\ 6 & -1 \\ 7 & -3 \\ 8 & -5 \\ 9 & -7 \\ 10 & -9 \\ 11 & -11 \\ 12 & -13 \\ 13 & -15 \\ 14 & -17 \\ 15 & -19 \\ 16 & -21 \\ 17 & -23 \\ 18 & -25 \\ 19 & -27 \\ \end{array}\]